MICQ(1) MICQ(1) mICQ−ICQ micq[-h][-c][-i<(locale)>][-b<>][-f<>][-v[]][-l<->] mICQICQ,.mICQ,,,,SMS,,,,,(UTF-8),,. −h|−−help . −c|−−nocolor . −i|−−i18n<(locale)> (locale). −b|−−basedir<> mICQ,micqrcfile. −f|−−config<> . −v|−−verbose[] ,. −l|−−logplace| -,-. mICQ,-,OS/2,Win32,BeOS-,Amiga,GNUHURDUnix.mICQ0.4.10,AIX, HPUX,Tru64,Amiga,BeOS-*.,,,. 16.1998. .(MatthewD.Smith)micq.1225.1998. 5.2001. ..,BSDmICQ-. 10.2002. 0.4.9GPL,8ICQ.(RdigerKuhlmann). - mICQ-micq-list@micq.org.http://www.micq.org/pipermail/micq-list/. HTTPhttp://www.micq.org/download. CVS-,: cvs−d:pserver:anonymous@micq.org:/var/lib/cvscomicq micqrc(5),micq(7) ˜ /.micqrc mICQ−. ˜ /.micq/ mICQ. /micqrc mICQ. SRYU mICQ 1 MICQ(1) MICQ(1) /i18n/*.i18n mICQ-. /history/ -. /files/. /share/micq/*.i18n mICQ. <micq@ruediger-kuhlmann.de>(). (JamesMorrison)<ja2morrison@student.math.uwaterloo.ca><jordi@sindominio.net>Debian /.:<lecicct@ptt.yu>. SRYU mICQ 2